# DALK: Dynamic Co-Augmentation of LLMs and KG to answer Alzheimer’s Disease Questions with Scientific Literature

### v2

@@ Abstract
Recent advancements in large language models (LLMs) have achieved promising performances across various applications. Nonetheless, the ongoing challenge of integrating long-tail knowledge continues to impede the seamless adoption of LLMs in specialized domains. In this work, we introduce DALK, a.k.a. Dynamic Co-Augmentation of LLMs and KG, to address this limitation and demonstrate its ability on studying Alzheimer’s Disease (AD), a specialized sub-field in biomedicine and a global health priority. With a synergized framework of LLM and KG mutually enhancing each other, we first leverage LLM to construct an evolving AD-specific knowledge graph (KG) sourced from AD-related scientific literature, and then we utilize a coarse-to-fine sampling method with a novel self-aware knowledge retrieval approach to select appropriate knowledge from the KG to augment LLM inference capabilities. The experimental results, conducted on our constructed AD question answering (ADQA) benchmark, underscore the efficacy of DALK. Additionally, we perform a series of detailed analyses that can offer valuable insights and guidelines for the emerging topic of mutually enhancing KG and LLM. We will release the code and data at https://github.com/David-Li0406/DALK.
@@ 1 Introduction
Nevertheless, directly applying these strategies in the context like AD would still suffer from several issues. First, Data Quality: As in many biomedical fields, scientific literature composes the largest publicly available corpus source in AD. Yet, the dense and information-overloaded nature of scientific literature, when combined with automatic retrieval methods, can lead to the retrieval of irrelevant and noisy information. Previous research has shown that noisy and irrelevant corpora can significantly undermine the performance of LLMs Yu et al. (2023); Chen et al. (2024); Wu et al. (2024). Second, Efficiency & Scale Issues: Being an critical field of research, the knowledge of AD is rapidly evolving with scientific advancements at a remarkable pace and scale. However, retraining a domain-specific LLM or updating certain knowledge in it demands substantial computational resources Hu et al. (2021); Ovadia et al. (2023); Zhang et al. (2024). This efficiency issue would also limit the sizes of domain-specific LLMs, consequently affecting their performances.
@@ 2 Related Work
KGs Miller (1995); Speer et al. (2017); Vrandečić and Krötzsch (2014) serve as structured representations of factual knowledge, typically expressed as (head, relation, tail) triples. Their structured, factual, and interpretable nature renders them excellent complements to parametric language models Pan et al. (2024). Recently, with the rise of large language models (LLMs), numerous studies have delved into exploring the synergy between LLMs and KGs for various purposes Pan et al. (2024); Tan et al. (2024). There are a lot of efforts in conducting knowledge graph construction Carta et al. (2023); Wadhwa et al. (2023); Han et al. (2023); Bi et al. (2024); Datta et al. (2024), completion Wei et al. (2023); Zhang et al. (2023b); Li et al. (2024) with the aid of LLMs. Conversely, other works aim to enhance LLMs by integrating knowledge sampled from KGs during both training Tang et al. (2023); Luo et al. (2024); Dernbach et al. (2024); Rangel et al. (2024) and inference Kim et al. (2023); Wen et al. (2023); Jiang et al. (2023); Sun et al. (2023a) times. Our work distinguishes itself by proposing a co-augmentation framework for LLMs and KGs, facilitating their mutual enhancement, and applying it to the domain of AD.
@@ 3 Our Methodology
Table 1: Detailed statistics about our augmented knowledge graph.
@@ 3.2.2 Self-aware Knowledge Retrieval
In our initial experiment, we noticed the coarse-grained knowledge sampled with the above-mentioned approaches still contained redundant and irrelevant information. This issue of noise is a common challenge encountered in automatically-constructed knowledge graphs Fang et al. (2021); Zhang et al. (2020); Li et al. (2022); Bi et al. (2024). Moreover, many recent works Yu et al. (2023); Li et al. (2023d); Chen et al. (2024); Wu et al. (2024) have demonstrated LLMs can indeed be influenced by such noisy information. To address this challenge, we borrow insights from the recent self-powered LLMs Wang et al. (2022); Pan et al. (2023); Li et al. (2023a); Yuan et al. (2024); Tong et al. (2024) and propose a self-aware knowledge retrieval method to leverage LLMs’ ranking capability Sun et al. (2023b); Ma et al. (2023) to filter out noisy information.
@@
We provide detailed examples in Appendix A and B to demonstrate the input and output in our DALK.
@@ 4.1 ADQA Benchmark
For performance evaluation, we consider four widely-used medical QA datasets spanning diverse biomedical topics Jin et al. (2021); Pal et al. (2022); Hendrycks et al. (2021); Peñas et al. (2013) and derive an AD-specific QA dataset from them. The four medical QA datasets are all multiple-choice based and include: 1) MedQA Jin et al. (2021) consisting of US Medical Licensing Examination (USMLE)-style questions, 2) MedMCQA Pal et al. (2022) containing medical school entrance exam questions from India, 3) MMLU Hendrycks et al. (2021) consisting of diverse biomedical and clinical questions from various sources, 4) QA4MRE Peñas et al. (2013) containing a subset of questions for AD derived from PubMed and Medline. In order to extract from the medical QA datasets a subset of samples related to AD for our evaluation, we referred to NIH’s Common Alzheimer’s and Related Dementias Research Ontology (CADRO)  [^1] https://iadrp.nia.nih.gov/about/cadro. Jointly developed by the National Institute on Aging and the Alzheimer’s Association, CADRO is a three-tiered classification system with eight main categories and a dozen sub-categories for AD and related dementia, and it contains common terminologies or keywords used in the field. We derived from the CADRO a list of AD-related keywords most relevant to the medical QA datasets: <Aging, Alzheimer, Amyloid beta, APOE, Dementia, Lipoprotein, Microglia>. Then, we searched against each medical QA dataset for matches with these keywords to find putative QA samples, then further asked GPT-3.5-turbo to judge for each putative sample whether the question is indeed related to AD or not. Finally, we filtered out a subset of such samples that are considered highly relevant to AD to conduct our evaluation (number of samples in each dataset is shown in Table 2). More details about ADQA can be found in Appendix C.
@@ 4.2 Experiment Settings
Both ChatDoctor Yunxiang et al. (2023) and Med-Alpaca Shu et al. (2023) are fine-tuned versions of LLaMA Touvron et al. on biomedical corpora. Compared with them, Meditron Chen et al. (2023) is built on LLaMA-2 Touvron et al. (2023) and extends its pretraining on a comprehensively curated medical corpus. BiomedGPT Zhang et al. (2023a) is also based on LLaMA-2 and pioneer as the first open-source and generalist visual language AI for diverse biomedical tasks. Biomistral Labrak et al. (2024) is an open-source LLM crafted specifically for the biomedical domain, optimized for efficiency through quantization and model merging techniques.
@@
ChatDoctor Yunxiang et al. (2023) 25.7 36.4 46.9 51.4 40.1
Med-Alpaca Shu et al. (2023) 41.4 42.8 44.9 57.1 46.5
BiomedGPT Zhang et al. (2023a) 38.8 41.9 48.9 42.6 43.1
Meditron Chen et al. (2023) 27.6 31.4 36.7 25.7 30.4
Biomistral Labrak et al. (2024) 44.7 49.5 53.1 68.6 54.0
@@
Table 2: Experiment results on our constructed ADQA benchmark. The sample size is indicated after each dataset name. The best results of each metric are in bold and the second-best results are underlined. The “AVG” column represents the average accuracy score on the four sub-dataset.
@@ 4.3 Main Result
Furthermore, we observe that the performance of biomedical-specific LLMs generally lags behind that of GPT-3.5-turbo. We attribute this discrepancy to the smaller size of these biomedical LLMs. While they may perform adequately in general medical contexts, they fall short in the AD scenario, which demands more domain-specific knowledge. In the case of GPT-3.5-turbo combined with various RAG methods, it is evident that most RAG methods enhance the models’ performance. Among them, GPT-3.5-turbo with Clinfo.ai yields the most significant improvement, boosting the accuracy score from 67.1 to 70.1 compared to vanilla GPT-3.5-turbo. However, it is important to note that the original Clinfo.ai necessitates access to the PubMed API, constituting an external resource. When we disable this access and solely utilize the same corpora as in DALK within the Clinfo.ai retrieval system, the improvement it brings becomes marginal and incomparable to our method. Due to the space limitation, we put more RAG results with different hyper-parameters in Appendix D.
@@ 5.1 Co-augmentation Analysis
Table 5: A case to show the effectiveness of DALK. The question is: “The area of the brain resistant to Neurofibrillary tangles of Alzheimer’s disease is: A. Visual association areas B. Entorhinal coex C. Temporal lobe D.Lateral geniculate body”
@@ 5.2 Hyper-parameter Analysis
In this section, we do a hyper-parameter analysis on the retrieval number $k$ of our self-aware retrieval module. We select a group of value for $k$ ([1,3,5,10,20,30]) and present the experiment results in Figure 4. We show the accuracy score on MedQA, MedMCQA, QA4MRE and AVG with different $k$. We find when $k$ is small, an increment to it can lead to a performance enhancement. After the best performance shows up, continually increasing the value of $k$ will cause a smooth decrease in the model accuracy score. This result indicates the knowledge ranked in the top positions is more helpful while the knowledge ranked behind is something not very useful, thus successfully validating the capability of LLMs to do a fine-grained knowledge reranking.
Moreover, we find the best $k$ value is correlated with the length of queries in each single dataset. For example, the best performance in MedQA (average query length is 107.4) shows up when $k=10$ while the best performance in MedMCQA and QA4MRE shows up when $k=5$ and $3$ respectively. This is consistent with our findings in Section 4.4 that a longer query corresponds to a larger and noisier sub-knowledge graph.
@@ 6 Conclusion
In this research, we begin by analyzing the main limitations of adopting the existing LLMs-based methods in AD-specific areas. To address these issues, we propose a novel approach in the merging of large language models and knowledge graphs in the context of Alzheimer’s Disease. Our team provides an innovative dynamic co-augmentation framework for the refinement of large language models and knowledge graphs. Initially, our approach extracts structural insights from the unstructured scientific literature, crafting a specialized knowledge graph for AD. Subsequently, we employ a coarse-to-fine sampling technique coupled with a unique self-aware knowledge retrieval strategy to pinpoint relevant information from the knowledge graph. The extensive evaluation conducted in our constructed ADQA benchmark showcases the effectiveness of our method and provides further hints into the synergy of LLMs and knowledge graph in the context of AD. In the future, we will do more exploration in adopting and benchmarking LLMs in the AD areas.
7 Limitations
In the development of our AD-KG, our primary focus lies in the exploration of two distinct methods for extracting relationships between associated entities. For entity recognition, we employ a strong PubTator annotator directly, without delving into the utilization of LLMs in this context. However, we have observed that LLMs also exhibit promising entity extraction capabilities in Section 3.2.1. We defer the refinement of methods for extracting entities for KG construction with LLMs to future works. Furthermore, a significant contribution of our work is the establishment of the ADQA benchmark. Nonetheless, the datasets utilized in constructing ADQA primarily consist of medical school exam questions, potentially exhibiting a domain gap from the scientific literature informing AD-KG. One potential remedy is leveraging PubmedQA Jin et al. (2019); however, it is hindered by limited data amount. In the future, we will keep gathering AD-related QA samples and expanding the size of our ADQA benchmark.
@@ 8 Ethics Statement
Appendix A Details of LLMs for KG
Table 7 and 8 present examples of our two KG construction methods respectively. For both methods, we adopt a select-or-generate prompt to instruct the LLM whether to choose a relation from hetionet Himmelstein et al. (2017), a well-built general medical KG, or generate a new one to describe the relationship between two entities. In the RE construction method, we also conduct a type matching (Table 9) for each entity from type name of PubTator to that of Hetionet and ask the LLM to choose from the relation set that corresponds to the two entities’ types (Table 10).

### v3

@@ Abstract
Recent advancements in large language models (LLMs) have achieved promising performances across various applications. Nonetheless, the ongoing challenge of integrating long-tail knowledge continues to impede the seamless adoption of LLMs in specialized domains. In this work, we introduce DALK, a.k.a. Dynamic Co-Augmentation of LLMs and KG, to address this limitation and demonstrate its ability on studying Alzheimer’s Disease (AD), a specialized sub-field in biomedicine and a global health priority. With a synergized framework of LLM and KG mutually enhancing each other, we first leverage LLM to construct an evolving AD-specific knowledge graph (KG) sourced from AD-related scientific literature, and then we utilize a coarse-to-fine sampling method with a novel self-aware knowledge retrieval approach to select appropriate knowledge from the KG to augment LLM inference capabilities. The experimental results, conducted on our constructed AD question answering (ADQA) benchmark, underscore the efficacy of DALK. Additionally, we perform a series of detailed analyses that can offer valuable insights and guidelines for the emerging topic of mutually enhancing KG and LLM.
@@ 1 Introduction
Nevertheless, directly applying these strategies in the context like AD would still suffer from several issues. First, Data Quality: Same as many biomedical fields, scientific literature composes the largest publicly available corpus source in AD. Yet, the dense and information-overloaded nature of scientific literature, when combined with raw text retrieval methods, can lead to the retrieval of irrelevant and noisy information. Previous research has shown that noisy and irrelevant corpora can significantly undermine the performance of LLMs Yu et al. (2023); Chen et al. (2024); Wu et al. (2024). Second, Efficiency & Scale Issues: Being an critical field of research, the knowledge of AD is rapidly evolving with scientific advancements at a remarkable pace and scale. However, retraining a domain-specific LLM or updating certain knowledge in it demands substantial computational resources Hu et al. (2021); Ovadia et al. (2023); Zhang et al. (2024). This efficiency issue would also limit the sizes of domain-specific LLMs, consequently affecting their performances.
@@ 2 Related Work
KGs Miller (1995); Speer et al. (2017); Vrandečić and Krötzsch (2014); Xiong et al. (2023, 2024) serve as structured representations of factual knowledge, typically expressed as (head, relation, tail) triples. Their structured, factual, and interpretable nature renders them excellent complements to parametric language models Pan et al. (2024); Zhu et al. (2023). Recently, with the rise of large language models (LLMs), numerous studies have delved into exploring the synergy between LLMs and KGs for various purposes Pan et al. (2024); Tan et al. (2024). There are a lot of efforts in conducting knowledge graph construction Carta et al. (2023); Wadhwa et al. (2023); Han et al. (2023); Bi et al. (2024); Datta et al. (2024), completion Wei et al. (2023); Zhang et al. (2023b); Li et al. (2024) with the aid of LLMs. Conversely, other works aim to enhance LLMs by integrating knowledge sampled from KGs during both training Tang et al. (2023); Luo et al. (2024); Dernbach et al. (2024); Rangel et al. (2024) and inference Kim et al. (2023); Wen et al. (2023); Jiang et al. (2023); Sun et al. (2023a) times. Our work distinguishes itself by proposing a co-augmentation framework for LLMs and KGs, facilitating their mutual enhancement, and applying it to the domain of AD.
@@ 3 Our Methodology
Table 1: Detailed statistics about AD-KG.
@@ 3.2.2 Self-aware Knowledge Retrieval
In our initial experiment, we noticed the coarse-grained knowledge sampled with the above-mentioned approaches still contained redundant and irrelevant information. This issue of noise is a common challenge encountered in automatically-constructed knowledge graphs Fang et al. (2021); Zhang et al. (2020); Li et al. (2022); Bi et al. (2024). Moreover, many recent works Yu et al. (2023); Li et al. (2023d); Chen et al. (2024); Wu et al. (2024) have demonstrated LLMs can indeed be influenced by such noisy information. To address this challenge, we borrow insights from the recent self-powered LLMs Wang et al. (2022); Pan et al. (2023); Li et al. (2023a); Yuan et al. (2024); Tong et al. (2024) and LLM-as-a-judge Zheng et al. (2023), proposing a self-aware knowledge retrieval method to leverage LLMs’ ranking capability Sun et al. (2023b); Ma et al. (2023) to filter out noisy information.
@@
We provide detailed examples in Appendix B and C to demonstrate the input and output in our DALK.
@@ 4.1 ADQA Benchmark
For performance evaluation, we consider four widely-used medical QA datasets spanning diverse biomedical topics Jin et al. (2021); Pal et al. (2022); Hendrycks et al. (2021); Peñas et al. (2013) and derive an AD-specific QA dataset from them. The four medical QA datasets are all multiple-choice based and include: 1) MedQA Jin et al. (2021) consisting of US Medical Licensing Examination (USMLE)-style questions, 2) MedMCQA Pal et al. (2022) containing medical school entrance exam questions from India, 3) MMLU Hendrycks et al. (2021) consisting of diverse biomedical and clinical questions from various sources, 4) QA4MRE Peñas et al. (2013) containing a subset of questions for AD derived from PubMed and Medline. In order to extract from the medical QA datasets a subset of samples related to AD for our evaluation, we referred to NIH’s Common Alzheimer’s and Related Dementias Research Ontology (CADRO)  [^1] https://iadrp.nia.nih.gov/about/cadro. Jointly developed by the National Institute on Aging and the Alzheimer’s Association, CADRO is a three-tiered classification system with eight main categories and a dozen sub-categories for AD and related dementia, and it contains common terminologies or keywords used in the field. We derived from the CADRO a list of AD-related keywords most relevant to the medical QA datasets: <Aging, Alzheimer, Amyloid beta, APOE, Dementia, Lipoprotein, Microglia>. Then, we searched against each medical QA dataset for matches with these keywords to find putative QA samples, then further asked GPT-3.5-turbo to judge for each putative sample whether the question is indeed related to AD or not. Finally, we filtered out a subset of such samples that are considered highly relevant to AD to conduct our evaluation (number of samples in each dataset is shown in Table 2). More details can be found in Appendix D.
@@ 4.2 Experiment Settings
Both ChatDoctor Yunxiang et al. (2023) and Med-Alpaca Shu et al. (2023) are fine-tuned versions of LLaMA Touvron et al. on biomedical corpora. Compared with them, Meditron Chen et al. (2023) is built on LLaMA-2 Touvron et al. (2023) and extends its pretraining on a comprehensively curated medical corpus. BiomedGPT Zhang et al. (2023a) is also based on LLaMA-2 and pioneer as the first open-source and generalist visual language AI for diverse biomedical tasks. Biomistral Labrak et al. (2024) is an LLM crafted specifically for the biomedical domain, optimized for efficiency through quantization and model merging techniques.
@@
ChatDoctor-7B Yunxiang et al. (2023) 25.7 36.4 46.9 51.4 40.1
Med-Alpaca-7B Shu et al. (2023) 41.4 42.8 44.9 57.1 46.5
BiomedGPT-7B Zhang et al. (2023a) 38.8 41.9 48.9 42.6 43.1
Meditron-7B Chen et al. (2023) 27.6 31.4 36.7 25.7 30.4
Biomistral-7B Labrak et al. (2024) 44.7 49.5 53.1 68.6 54.0
Meditron-70B 50.0 44.8 79.6 51.4 56.4
ClinicalCamel-70B Toma et al. (2023) 50.0 64.3 83.7 68.6 66.7
@@
Table 2: Experiment results on our constructed ADQA benchmark. The best results of each metric are in bold and the second-best results are underlined. The “AVG” column represents the average accuracy score on the four sub-dataset.
@@ 4.3 Main Result
Furthermore, we observe that the performance of biomedical-specific LLMs generally lags behind that of GPT-3.5-turbo. We attribute this discrepancy to the smaller size of these biomedical LLMs. While they may perform adequately in general medical contexts, they fall short in the AD scenario, which demands more domain-specific knowledge. In the case of GPT-3.5-turbo combined with various RAG methods, it is evident that most RAG methods enhance the models’ performance. Among them, GPT-3.5-turbo with Clinfo.ai yields the most significant improvement, boosting the accuracy score from 67.1 to 70.1 compared to vanilla GPT-3.5-turbo. However, it is important to note that the original Clinfo.ai necessitates access to the PubMed API, constituting an external resource. When we disable this access and solely utilize the same corpora as in DALK within the Clinfo.ai retrieval system, the improvement it brings becomes marginal and incomparable to our method. Due to the space limitation, we put more RAG results with different hyper-parameters in Appendix E.
@@ 5.1 Co-augmentation Analysis
Table 5: An example for the case study. The question is: “The area of the brain resistant to Neurofibrillary tangles of Alzheimer’s disease is: A. Visual association areas B. Entorhinal coex C. Temporal lobe D.Lateral geniculate body”
@@ 5.2 Hyper-parameter Analysis
In this section, we do a hyper-parameter analysis on the retrieval number $k$ of our self-aware retrieval module. We select a group of value for $k$ ([1,3,5,10,20,30]) and present the experiment results in Figure 4. We show the accuracy score on MedQA, MedMCQA, QA4MRE and AVG with different $k$. We find when $k$ is small, an increment to it can lead to a performance enhancement. After the best performance shows up, continually increasing the value of $k$ will cause a smooth decrease in the model accuracy score. This result indicates the knowledge ranked in the top positions is more helpful while the knowledge ranked behind is not very useful, thus successfully validating the capability of LLMs to do fine-grained knowledge reranking.
Moreover, we find the best $k$ value is correlated with the averaged query length. For example, the best performance in MedQA (average query length is 107.4) appears when $k=10$ while the best performance in MedMCQA and QA4MRE appears when $k=5$ and $3$ respectively. This aligns with our previous finding that a longer query corresponds to a larger and noisier sub-knowledge graph.
@@ 5.4 Case Study
5.5 Experiment Results with GPT-4
MedMCQA MMLU
GPT-4-turbo 88.1 81.3
GPT-4-turbo w/ DALK 89.2 91.8
Table 7: Experiment Result in GPT-4.
To evaluate our method in the SOTA LLM, we select two subsets, MedMCQA and MMLU and conduct the experiment with GPT-4-turbo. As the result shown in the Table 7, we find in both subsets, DALK achieves significant performance enhancement. Another thing to note is in the MMLU dataset, vanilla GPT-4-turbo (81.3) is even worse than GPT-3.5-turbo (83.6). We will conduct more experiments in larger models in the future to thoroughly explore our proposed method.
5.6 Generalization of DALK
GPT-4 Score
GPT-3.5-turbo 0.61
GPT-3.5-turbo w/ DALK 0.66
Table 8: Experiment Result in COVID-QA.
To demonstrate the robustness and generalization of our DALK, We conduct a further evaluation with COVID-QA, an open-ended QA dataset focused on COVID-19. For efficiency, We use a small set of literature abstracts from the CORD-19 dataset as the raw corpus to build KG. We conduct the experiment using GPT-3.5-turbo from OpenAI. For evaluation, we follow some previous work and prompt GPT-4 to score each generated answer. As the result shown in Table 8, DALK also significantly improves model performance in this open-ended QA dataset focuses on COVID.
@@ 6 Conclusion
In this research, we begin by analyzing the main limitations of adopting the existing LLMs-based methods in AD-specific areas. To address these issues, we propose a novel approach in the merging of large language models and knowledge graphs in the context of Alzheimer’s Disease. We provide an innovative dynamic co-augmentation framework for the refinement of large language models and knowledge graphs. Initially, our approach extracts structural insights from the unstructured scientific literature, crafting a specialized knowledge graph for AD. Subsequently, we employ a coarse-to-fine sampling technique coupled with a unique self-aware knowledge retrieval strategy to pinpoint relevant information from the knowledge graph. The extensive evaluation conducted in our constructed ADQA benchmark shows the effectiveness of our method and provides further hints into the synergy of LLMs and KG in the context of AD.
7 Limitations and Future Work
In the development of our AD-KG, our primary focus lies in the exploration of two distinct methods for extracting relationships between associated entities. For entity recognition, we employ a strong PubTator annotator directly, without delving into the utilization of LLMs in this context. However, we have observed that LLMs also exhibit promising entity extraction capabilities in Section 3.2.1. We defer the refinement of methods for extracting entities for KG construction with LLMs to future works. Furthermore, a significant contribution of our work is the establishment of the ADQA benchmark. Nonetheless, the datasets utilized in constructing ADQA primarily consist of medical school exam questions, potentially exhibiting a domain gap from the scientific literature informing AD-KG. One potential remedy is leveraging PubmedQA Jin et al. (2019); however, it is hindered by limited data amount. In the future, we will keep gathering AD-related QA samples and expanding the size of our ADQA benchmark. In the future, we will do more exploration in adopting and benchmarking LLMs in the AD areas.
@@ 8 Ethics Statement
Acknowledgments
This work was supported in part by the NIH grants U01 AG066833, U01 AG068057, R01 AG071470, and P30 AG073105. The material in this presentation is also based upon work supported by the U.S. Department of Homeland Security under Grant Award Number, 17STQAC00001-08-00 and the U.S. Office of Naval Research (ONR) under grant N00014-21-1-4002.
Appendix A AD-KG Evaluation
We conduct a manual evaluation of our AD-KG to guarantee its quality. We randomly sample 100 triples from AD-KG constructed with generative and RE methods and ask the annotator to check whether the fact in each triple is valid. The annotator is encouraged to find evidence from the original abstract corpus or by searching the web. The results are presented in Table 9. We found that while both AD-KGs achieve acceptable accuracy, the RE construction method produces an AD-KG of higher quality. This further validates our conclusion in Section 4.5 regarding the trade-off between coverage and accuracy.
Accuracy
AD-KG w/ Generative Construction 91%
AD-KG w/ RE Construction 83%
Table 9: Manual evaluation result on AD-KG.
Appendix B Details of LLMs for KG
Table 10 and 11 present examples of our two KG construction methods respectively. For both methods, we adopt a select-or-generate prompt to instruct the LLM whether to choose a relation from hetionet Himmelstein et al. (2017), a well-built general medical KG, or generate a new one to describe the relationship between two entities. In the RE construction method, we also conduct a type matching (Table 12) for each entity from type name of PubTator to that of Hetionet and ask the LLM to choose from the relation set that corresponds to the two entities’ types (Table 13).